OpenStudy (madmerc):

Find the solution of 3 sqrt x+5 = -9 and determine if it is an extraneous solution x = 4 ; not extraneous x = 4 ; extraneous x = 0 ; extraneous x = 0 ; not extraneous

OpenStudy (madmerc):

\[3\sqrt{x + 5}=-9\] is what the question looks like

OpenStudy (madmerc):

@freckles please help I don't understand what this is

OpenStudy (baru):

well we can solve this equation..take 3 to rhs, square both sides blah blah... but look at the equation carefully, we cannot get a negative value for square root. although we can solve for x and find a number, it is 'extraneous'

OpenStudy (madmerc):

? i don't understand what that means. the solution is extraneous? but what does x equal?

OpenStudy (baru):

ok...lets first solve for x \(\\\sqrt{x+5} = -3 \\square~ both ~sides\\ x+5=9\\x=4\)

OpenStudy (baru):

followed?

OpenStudy (madmerc):

yeah okay ! so it's the second one, x = 4 extraneous

OpenStudy (baru):

do you understand why its extraneous?

OpenStudy (madmerc):

something to do with the solution right like its not proven right

OpenStudy (baru):

yep, so we have \( \sqrt{x+5}=-3\) do you know any number whose square root is negative?

OpenStudy (madmerc):

no...

OpenStudy (baru):

because there isnt... this equation actually has no solution. but we have somehow solved for x using correct mathematical operations. that is why the solution is called 'extraneous'